if a^3b = abc = 180, where a,b,c are integers, then value of c is
Answer by Alan3354(69443)

a^3b = abc = 180, where a,b,c are integers
------------
a^2b = bc
a^2 = c
----------------
1 number has to be a multiple of 5, and it can't be a.
--> b is a multiple of 5.
a = 1
b = 180
c = 1